You’ve observed the following returns on Crash-n-Burn Computer’s stock over the past five years: 14 percent, –14 percent, 16 percent, 26 percent, and 10 percent. Suppose the average inflation rate over this period was 3.5 percent and the average T-bill rate over the period was 4 percent.

a.

What was the average real risk-free rate over this time period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What was the average real risk premium? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
a. Average real risk-free rate %
b. Average real risk premium %
Answer #1

Average real risk free rate= (1 +R) = (1 +1)(1 +h) R = 0.48% (1+4%)/(1.035)-1 Average real risk free rate = 0.48% Average ret